OpenStudy (agentnao):

A crew practices rowing up a river for a fixed distance and then returning to their starting point. The river has a current of 3km/h. The crew can row 15km/p in still water and if it takes them 25 minutes to make a round trip, how far upstream did they row?

OpenStudy (agentnao):

So what do you think? Do you know how to do this type of problem?

OpenStudy (anonymous):

I am here, give me a minute I will try this first on my copy.

OpenStudy (agentnao):

Alright.

OpenStudy (anonymous):

Okay Lets take that fixed distance as d. |----------------------------------------------------------------------------------------------| Starting Final point point The distance is d. km Now while going downstream or rowing along the river Makes the velocity 15+3=18km/h And while against the river means upstream 15 -3 =12 So time \(t_1\) while going upstream is \(\frac{d}{12}\) and downstream \(t_2\) is \(\frac{d}{18}\) Now \(t_1\) + \(t_2\) = \(\frac{25}{60}\) Divided by 60 to convert into hour from minutes. \[d(\frac{1}{12} + \frac{1}{18}) = \frac{25}{60}\]

OpenStudy (anonymous):

d is 3 km.

OpenStudy (anonymous):

And that must be your answer : )

OpenStudy (agentnao):

Yes! 3 is indeed the answer :) But give me a second to understand what you've done.

OpenStudy (anonymous):

Sure

OpenStudy (agentnao):

Okay, so this is a silly question, but is the 3 km referring to the distance traveled UPstream or the distance traveled totaled?

OpenStudy (agentnao):

Meaning, from starting point to ending point and back.

OpenStudy (anonymous):

No just upstream

OpenStudy (agentnao):

How did you figure that out though? How did you know that was just upstream? Does that mean the total distance would be 6?

OpenStudy (anonymous):

Yes

OpenStudy (agentnao):

How did you get 3 then? The d(1/18+1/12) = 25/60 should have told you the total distance, shouldn't it?

OpenStudy (anonymous):

No try it your self

OpenStudy (anonymous):

d is the distance for one trip

OpenStudy (agentnao):

I did. The problem stated that it took the crew 25 minutes to do the round trip, meaning the distance it took them to get from start to finish, so how does combining the different rates give only one leg of it?

OpenStudy (agentnao):

Oh. Ohhh. I think I understand what you did now. Okay, I see.

OpenStudy (anonymous):

ah Naomi the distance is same from start to final and to make round trip we travel twice the distance once from start to final and then from final to start.

OpenStudy (agentnao):

Yes, I see that now. The next part of the question says that the crew shortened the round trip time to 23 minutes. What is their still water speed now. Will I still use the d=rt formula for this?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

make that 2d instead of d.